# Multiway Alignment of Political Attitudes

### v1

@@ Abstract
The related concepts of partisan belief systems, issue alignment, and partisan sorting are central to our understanding of politics. These phenomena have been studied using measures of alignment between pairs of topics, or how much individuals’ attitudes toward a topic reveal about their attitudes toward another topic. We introduce a higher-order measure that extends the assessment of alignment beyond pairs of topics by quantifying the amount of information individuals’ opinions on one topic reveal about a set of topics simultaneously. Our multiway alignment measure indicates how much individuals’ opinions on multiple topics align into a single ideological divide. Applying this approach to legislative voting behavior reveals that parliamentary systems typically exhibit similar multiway alignment characteristics, but can change in response to shifting intergroup dynamics. In American National Election Studies surveys, our approach reveals a growing significance of party identification together with a consistent rise in multiway alignment over time. Similarly, the growing multiway alignment among topical issues in Finnish online discussions suggests a trend towards a more ideologically driven political landscape. Our case studies demonstrate that the multiway alignment measure is a versatile tool for understanding societal polarization and partisan belief systems across diverse domains.
@@ 1 Introduction
When examining ideological cleavages across multiple topics and multidimensional constraints, researchers can use multivariate measures that capture global characteristics of joint distributions (e.g., Rosas et al., 2019), such as the extent of information shared among variables and the level of interdependence, through the entropies of the stochastic variables. We document these in E. Some of these measures, however, present applicational challenges in that they do not have directly correspondent bivariate analyses. Moreover, the use of unnormalized measures complicates the interpretability of the scores and comparisons across different datasets.
@@
Our approach to multiway alignment analysis facilitates understanding the interplay between various divides and beliefs. We illustrate this using multiple cases that vary by their political and geographical context. Our analysis of data from the American National Election Studies (ANES; 2021) reveals the increasingly substantial influence of party identification in shaping opinion dynamics among the U.S. public. In Finnish online political discourse, we observe a pronounced increase in higher-order alignment across issues, signaling a transformation towards a more ideologically driven political environment. Our examination of the Finnish Parliament before and during the COVID pandemic highlights a shift toward considerably lower levels of higher-order alignment in response to the challenges posed by the crisis, a well-recognized phenomenon that pairwise analysis alone does not capture. Finally, we observe elite polarization from voting behavior in the U.S. House, where a multitude of policy issues manifests high alignment. To facilitate further exploration of our findings and application to other systems, we provide the Python package multiway_alignment (H), which lets researchers apply our multiway alignment methodology.
The paper proceeds as follows. Section 2 establishes the theoretical framework that extends the concept of alignment beyond the pairwise to effectively model higher-order interdependencies. Section 3 details our approach to quantifying multiway alignment. In Section 4, we demonstrate the broad applicability of our measure to a variety of social and political systems and we discuss our findings. Finally, Section 5 summarizes our contributions and outlines future research directions.
@@ 2 Importance of Multiway Alignment
We propose a multiway alignment measure that gauges the extent to which individuals’ attitudes toward different topics collapse into a single dimension. This concept goes beyond examining alignment over pairs of topics and instead investigates the alignment of opinions over an array of topics and the extent to which consensus among individuals transcends the specific subjects being discussed. To do so, we define the idea of consensus partition to capture the multidimensional nature of social systems. Specifically, a consensus partition refers to a division of individuals into groups (i.e., a consensus group) based on how they share opinions across multiple topics. As such, these partitions comprise groups of individuals who hold the same constellation of positions across a set of topics. We further formalize this concept in Section 3.
@@ 3 Measuring Multiway Alignment
In this section, we formally develop our multiway alignment measure. In Section 3.1, we formalize the idea of consensus partitions and its significance in investigating the extent to which opinion groups across multiple topics are either aligned or cross-cutting. In Section 3.2, we produce a list of desired properties for a multiway alignment measure and present our formulation for it.
@@ 3.2 Multiway Alignment
We start building our multiway aligned measure by listing desirable properties, and use this list to construct the measure:
@@
    The measure should be able to describe higher-order phenomena while still having an intuitive interpretation and correspondence with the relevant theoretical concepts of consensus and opinion groups we developed above.
@@
  4. 4.
    The measure should be designed to avoid overfitting. It should clearly differentiate between true alignment and alignment that can be attributed to random chance, possibly by allowing for statistical significance test or a baseline comparison.
@@
Given the interpretation of the consensus partition presented in Section 3.1, we establish a quantitative measure of multiway alignment across a set of $k$ topics based on how similar the opinion partition of each topic is to the consensus partition of the $k-1$ remaining topics. In particular, for a given list of topic $T_{1},\ldots,T_{k}$, we can quantify their multiway alignment by averaging a chosen similarity measure S between each topic and the consensus partition defined by the remaining topics:
If the similarity is based on mutual information, i.e., $S=MI$, the measure describes the average amount of information knowing someone’s opinion in one topic gives about their position in all of the other topics. As shown in D, $A_{\text{MI}}$ can be interpreted as the total correlation, adjusted by subtracting contributions from lower-order alignments.
@@
As long as the chosen S is symmetric, $A_{S}$ reduces to the original pairwise measure $S$ for $k=2$, since the consensus partition $C(T_{j})$ is defined by topic $T_{j}$ alone and therefore matches opinions in $T_{j}$ itself. By the same reasoning above, the formula in Equation 2 simplifies to exactly the measure used in Chen et al. (2021) for $k=2$ when S = NMI. Yet, the formulation in Equation 2 allows us to quantify also alignment across multiple topics, as the mutual information is now computed between each topic and the consensus partition defined by the remaining topics in analysis.
Property 3
@@
By definition, normalized mutual information (Vinh et al., 2009) takes on values between 00 and $1$, where 00 indicates independence and $1$ signifies complete agreement between the two labeling assignments that are compared. The $A_{NMI}$ score naturally inherits those properties, and therefore allows for similar interpretation. NMI, however, does not correct for agreement solely due to chance and thus may yield inflated scores, particularly when comparing groups with significant size imbalances (Newman et al., 2020; Salloum et al., 2022). This limitation also applies to the multiway alignment score $A_{NMI}$. We mitigate this issue by using the adjusted mutual information (AMI) (Vinh et al., 2009) instead, and all our numerical results will be presented in terms of $A_{AMI}$.
Property 4
We validate the statistical significance of our multiway alignment findings by introducing a null model for comparison (F). The null model provides the expected outcome $\langle A_{\text{null}}\rangle$ when the multiway alignment is a consequence of random chance and allows us to establish confidence intervals that account for fluctuations in the value of $A$ due to the finite number of individuals in the system.
A multiway alignment score defined as in Equation 2 with $S=AMI$ satisfies the four properties we stated above. Additionally, we define the maximal alignment curve (C) to describe the evolution of alignment as an increasing number of topics are considered. The maximal alignment curve collects the maximum alignment scores reached by the system at any possible combinations of $k$ topics. By considering the sets of topics exhibiting the highest alignment, the shape of the maximal alignment curve summarizes the overall alignment of the system, and the area under this curve can be used as a global summary statistic, describing the higher-order alignment of the system.
4 Applications
Figure 3: The shaded areas in Panel (a) show the range of values obtained for the multiway alignment across 43 different systems, while the lines represent the maximal alignment curves for selected periods and systems. Panel (b) and (c) show the spectrum of multiway alignment in Finnish Parliament over the parliamentary years 2018/2019 and 2020/2021 respectively. Panel (d) shows the spectrum of the multiway alignment obtained from the responses to ANES surveys in 2020. The spectrum in Panel (e) is obtained from the retweet networks before Finnish elections 2023. Panel (f) considers the votes in the U.S. House in 2014.
In this section, we illustrate higher-order alignment patterns revealed by our multiway alignment measure in case studies that cross national, institutional – public opinion surveys, political chambers, and social media platforms – and temporal contexts. Specifically, we analyzed Finnish Parliament votes from 1996–2021, ANES surveys collected every four years from 2004–2020, pre-election discussions on Twitter in 2019 and 2023, and U.S. House activities from 2004–2013. Figure 3 summarizes all results from these systems and years. The first panel, Figure 3a, shows with the shaded areas the range of multiway alignment values obtained from the 43 systems we analyzed, and examples of the maximal alignment curve. The remaining panels show the full spectrum of multiway alignment corresponding to each maximal alignment curve in Figure 3a. The spectrum includes all scores obtained when considering sets of topics of size two and above, and highlights values that are significant at $\alpha=5\%$.
Despite the distinct nature of each system, the multiway alignment spectrum observed from our datasets generally tends towards a funnel shape (Figure 3b-e-f and corresponding maximal alignment curves in Figure 3a for Finnish Parliament 2018–2019, Finnish Twitter 2023 and U.S. House 2014). Yet, the shape of the spectrum can vary substantially: the difference in Figure 3b-c and in the maximal alignment curves for Finnish Parliament 2018–2019 and 2020–2021 (Figure 3a) reflect a well-recognized impact of the pandemic, that we further discuss in Section 4.1.
Figures 3b-c (Finnish Parliament) and Figure 3f (U.S. House) demonstrate how parliamentary and governmental chambers often exhibit high multiway alignment due the collaborative nature of those legislative assemblies, where alignment reflects teamsmanship and elite polarization. Conversely, the level of higher-order alignment among the general public is lower, and further decreases when the sample of the population is random. This is evident from Figure 3a, where the values obtained when measuring alignment from ANES surveys are clearly the lowest among the four systems. Still, as discussed in Section 4.2 and Section 4.3, we observe an increasing trend in alignment also in the general population.
4.1 Multiway Alignment in Finnish Parliament Before and During COVID Pandemic
The analysis of multiway alignment within the Finnish Parliament offers insights into the shifting political dynamics before and during the COVID-19 pandemic, and reveals a known phenomenon that purely pairwise analysis would miss. The line chart in Figure 4 shows the evolution over time of the overall alignment, summarized through the area under the maximal alignment curve. The two points highlighted in the line chart show how the emergence of the pandemic and the subsequent declaration of a state of emergency led to a sharp decrease in alignment in 2020–2021.
@@ 4.1 Multiway Alignment in Finnish Parliament Before and During COVID Pandemic
A comparison of the multiway alignment spectrum for the 2019–2020 and 2020–2021 parliamentary periods is shown in the main panel in Figure 4. Before the onset of the pandemic, the spectrum of the multiway alignment within the Finnish Parliament displayed a funnel shape that we commonly observe in legislative systems, where the alignment across different topics reflects the dynamics of policymaking and political discourse. The funnel shape of multiway alignment indicates a gradual decrease in alignment as the number of topics in the analysis increases. Nevertheless, the values can remain at values above zero also at higher orders (e.g., $k=9$ and $k=10$ in 2019–2020), demonstrating alignment that could originate from adherence to party lines, whereby members from the same party (or coalition) vote in a similar way on a wide range of issues. [^1] The same effect is also visible in Figure 3f, which shows one instance of the alignment spectrum for U.S. House (Appendix G). Being a two-party system, the U.S. witnesses a much more pronounced effect of elite polarization on the spectrum of multiway alignment, and the multiway alignment score remains at moderate values of around 0.3 even when considering all the topics discussed over the year ($k=18$).
Notably, during the COVID-19 pandemic, especially early on, Finnish parliamentary politics exhibited unconventionally high levels of cooperation between the government and opposition coalitions (Lehtonen and Ylä-Anttila, 2024). This was evident in how both sides repeatedly appealed for broad cooperation, reciprocally praise each other for implemented COVID responses, and even in how the opposition withdrew their interpellation on the government’s immigration policy (Eduskunta.fi, 2020). Despite this drastic shift from conventional politics, the overall patterns gleaned from a pairwise alignment analysis exhibit no visible difference across the two periods. The only hint of a reorientation of priorities and focal points within parliamentary deliberations is a shift in the most aligned pair of topics, from defense and health in 2019–-2020 to transportation and health in 2020-–2021.
The changes brought by the pandemic become glaringly evident only when analyzing the patterns of alignment at higher orders: the sharp drop in multiway alignment following $k=7$ and the manifest shift in the distribution of the scores already at $k=3$ usher in a period of exceptional policy discussions. At $k=3$, the spectrum for 2020–2021 shows two clear clusters of topics, one retaining alignment levels comparable to the previous period, and the other exhibiting markedly low alignment. The same clustering effect becomes increasingly distinct at higher orders ($k=4,5,6$). Overall, the spectrum of alignment in 2020–2021 clearly shows six topics – transportation, culture, defense, economics, foreign affairs, and health – following similar alignment patterns as in 2019-2020. The remaining five topics – climate, education, immigration, social issues, and unemployment – become much less aligned in 2020–2021, bringing the multiway alignment score at $k=10$ and $k=11$ towards zero.
Looking more deeply into these patterns, we see that they reflect the collaborative dynamics explicitly adopted to face the crisis. On one hand, the sudden prominence of topics related to transportation and their alignment with health and economics suggests a focused effort within the Finnish Parliament to navigate the circumstances imposed by the pandemic through initiatives related to free movement restrictions, economic stimulus packages, quarantine and isolation protocols, and wide vaccination efforts (YLE, 2020a, b). On the other hand, topics such as climate and immigration could have been perceived as less immediately relevant or lower in priority compared to the urgent challenges posed by the crisis. Overall, our findings quantitatively support the observation that, throughout the pandemic, parliamentary discussions diverged from traditional political dynamics, and that the opposition adopted a collaborative approach, refraining from strong criticism of COVID policies proposed by the government (Niemikari and Raunio, 2022).
4.2 Evolution of Alignment in Finnish Twitter Before Elections
4.3 Multiway Alignment and Party Preferences: Higher-Order Partisan Sorting of the U.S. Public
@@ 4.2 Evolution of Alignment in Finnish Twitter Before Elections
Finally, we study multiway alignment (both issue alignment and partisan alignment) in U.S. public opinion using the ANES Time Series data from 2004–2020. First, we show in Figure 6 the full spectrum of the multiway alignment computed from the 2004 and 2020 data, which reveals a shifting salience of relationships over this temporal span among eight topics (seven issues and partisan identification). In particular, Figure 6 shows a transition from the predominant alignment of stances concerning abortion and LGBT rights in 2004 to a more pronounced alignment between gun access and immigration in 2020. Notably, the alignment score pertaining to the latter pair surpasses the three highest scores observed in 2004, signifying the existence of a stronger 2-way constraint in 2020. Beyond this pairwise analysis, Figure 6 also shows an increase over time in the overall multiway alignment across all eight topics, indicating intensified interdependencies among the analyzed stances and portraying a more polarized landscape of the contemporary American discourse.
@@
The analysis of 3-tuples and higher illustrates the growing importance of partisanship in U.S. politics. Whereas in 2004, positions toward immigration played a central role within triadic-and-higher relationships, in 2020, party preferences became one of the most prominent topics, marking a shift towards sociopolitical affiliations as key alignment determinants. In this respect, our measure of multiway alignment allows us to investigate research questions that study partisanship across a multitude of issues. Specifically, we are now able to quantify the strength of the relationship between party preference and a bundle of multiple topics.

### v2

@@ Abstract
The related concepts of partisan belief systems, issue alignment, and partisan sorting are central to our understanding of politics. These phenomena have been studied using measures of alignment between pairs of topics, or how much individuals’ attitudes toward a topic reveal about their attitudes toward another topic. We introduce a higher-order measure that extends the assessment of alignment beyond pairs of topics by quantifying the amount of information individuals’ opinions on one topic reveal about a set of topics simultaneously. Applying this approach to legislative voting behavior shows that parliamentary systems typically exhibit similar multiway alignment characteristics, but can change in response to shifting intergroup dynamics. In American National Election Studies surveys, our approach reveals a growing significance of party identification together with a consistent rise in multiway alignment over time.
@@ 1 Introduction
A common way to measure the degree of agreement among multiple individuals and across multiple issues is Kendall’s W (Gibbons and Kendall, 1990), which quantifies the agreement across different rankings of the same set of objects. Kendall’s W, however, is a linear function of the averages of the pairwise Spearman’s rank correlation coefficients between every possible pair (Gibbons and Kendall, 1990). As we show in the following section, the average of pairwise scores does not appropriately describe systems at higher orders. Alternatively, when examining ideological cleavages across multiple topics and multidimensional constraints, researchers can use multivariate measures that capture global characteristics of joint distributions (e.g., Rosas et al., 2019), such as the extent of information shared among variables and the level of interdependence, through the entropies of the stochastic variables. We document these in Appendix E. Some of these measures, however, present applicational challenges in that they do not have directly correspondent bivariate analyses. Moreover, the use of unnormalized measures complicates the interpretability of the scores and comparisons across different datasets.
@@
Our approach to multiway alignment analysis facilitates understanding the interplay between various divides and beliefs. Our analysis of data from the American National Election Studies (ANES; 2021) reveals the increasingly substantial influence of party identification in shaping opinion dynamics among the U.S. public. To facilitate further exploration of our findings and application to other systems, we provide the Python package multiway_alignment (see Appendix J), which lets researchers apply our multiway alignment methodology.
The paper proceeds as follows. First, we establish the theoretical framework that extends the concept of alignment beyond the pairwise to effectively model higher-order interdependencies. Then, we detail our approach to quantifying multiway alignment. We proceed to demonstrate the broad applicability of our measure to a variety of social and political systems and discuss general patterns of higher-order alignment that may be observed in real-world systems. We then discuss our findings from the analysis of data from the American National Election Studies (ANES; 2021). Finally, the last section summarizes our contributions and outlines future research directions.
@@ 2 Importance of Multiway Alignment
We propose a multiway alignment measure that gauges the extent to which individuals’ attitudes toward different topics collapse into a single divide. This concept goes beyond examining alignment over pairs of topics and instead investigates the alignment of opinions over an array of topics and the extent to which consensus among individuals transcends the specific subjects being discussed. To do so, we define the idea of consensus partition to capture the multidimensional nature of social systems. Specifically, a consensus partition refers to a division of individuals into groups (i.e., a consensus group) based on how they share opinions across multiple topics. As such, these partitions comprise groups of individuals who hold the same constellation of positions across a set of topics. We further formalize this concept in Section 3.
@@ 3 Measuring Multiway Alignment
In this section, we formally develop our multiway alignment measure. In Section 3.1, we formalize the idea of consensus partitions and its significance in investigating the extent to which opinion groups across multiple topics are either aligned or cross-cutting. In Section 3.2, we present our formulation for a multiway alignment measure and discuss its desired properties.
@@ 3.2 Multiway Alignment
Given the interpretation of the consensus partition presented in the previous Section, we establish a quantitative measure of multiway alignment across a set of $k$ topics based on how similar the opinion partition of each topic is to the consensus partition of the $k-1$ remaining topics. In particular, for a given list of topics $T_{1},\ldots,T_{k}$, we can quantify their multiway alignment by averaging a chosen similarity measure S between each topic and the consensus partition defined by the remaining topics:
[…]
The measure defined in Equation 2 is able to describe higher-order phenomena while still having an intuitive interpretation and correspondence with the relevant theoretical concepts of consensus and opinion groups we developed above. Even though the definition does not restrict the choice of the similarity measure S, we seek to choose S in a way that guarantees certain desirable properties for our measure:
@@
As long as the chosen S is symmetric, $A_{\text{S}}$ reduces to the original pairwise measure S for $k=2$, since the consensus partition $C(T_{j})$ is defined by topic $T_{j}$ alone and therefore matches opinions in $T_{j}$ itself. By the same reasoning above, the formula in Equation 2 simplifies to exactly the measure used in Chen et al. (2021) for $k=2$ when S = NMI. Yet, the formulation in Equation 2 allows us to quantify also alignment across multiple topics, as the mutual information is now computed between each topic and the consensus partition defined by the remaining topics in analysis.
@@
The normalization of mutual information is achieved by scaling MI by one of its upperbounds: for example, the maximum of the entropies, the geometric average of the entropies, or the arithmetic mean of the entropies (Vinh et al., 2009). As a result, normalized mutual information takes on values between 00 and $1$, where 00 indicates independence and $1$ signifies complete agreement between the two labeling assignments that are compared. The $A_{\text{NMI}}$ score naturally inherits those properties, and therefore allows for similar interpretation. NMI, however, does not correct for agreement solely due to chance and thus may yield inflated scores, particularly when comparing groups with significant size imbalances (Newman et al., 2020; Salloum et al., 2022). This limitation also applies to the multiway alignment score $A_{\text{NMI}}$. We mitigate this issue by using the adjusted mutual information (AMI) (Vinh et al., 2009) instead, and all our numerical results will be presented in terms of $A_{\text{AMI}}$. Specifically,
$\text{AMI}(U,V)=\frac{\text{MI}(U,V)-\mathbb{E}\{\text{MI}(C)|a_{1},\dots,a_{m}\}}{\text{max}\{\text{MI}(U,V)\}-\mathbb{E}\{\text{MI}(C)|a_{1},\dots,a_{m}\}}\,,$
where $\text{MI}(U,V)$ is the mutual information between $U$ and $V$, $\text{max}\{\text{MI}(U,V)\}$ is one of the upperbounds of MI that can be used to normalize MI, $a_{1},\dots,a_{m}$ are the marginals from the contingency table $C$ between $U$ and $V$ so that the expected mutual information $\mathbb{E}\{\text{MI}(C)\mid a_{1},\dots,a_{m}\}$ is conditioned on the overall distributions of the cluster sizes (Vinh et al., 2009).
[…]
4 Alignment Spectrum
In this section, we introduce the concept of the alignment spectrum and explain how it can be used to understand higher-order alignment patterns across multiple topics. The alignment spectrum emerges from the calculation of all alignment scores across all combinations of topics. Essentially, it illustrates how political attitudes relate to each other as we expand the set of topics being considered together. In fact, within a system of $k$ topics, we can calculate several alignment scores – specifically, pairwise alignments between each pair of topics, as well as all higher-order alignment scores that consider order $3,\ldots,k$. For example, in the case of four topics, the alignment spectrum will consist of the scores derived from pairs of topics, triplets of topics, and the full set of four topics. This results in a total of seven alignment scores that reflect how individuals’ opinions relate across different subsets of the four topics. By examining these scores, we can identify patterns of alignment and study how alignment evolves as we increase the number of topics.
We show an example of an alignment spectrum in Figure 3a, which illustrates the structure of online discussions around four topics (climate, economic policy, education, and immigration) in the Finnish Twittersphere leading up to the 2023 Finnish elections. The $x$-axis represents the order of alignment: 2 for pairs of topics, 3 for triplets, and 4 for the full set of topics. Each point on the plot corresponds to a specific subset of topics in the respective order, and the $y$-axis displays the alignment values for each of those subsets. The alignment spectrum summarizes different alignments in the system and shows how the patterns of alignment observed between pairs of topics extend to higher orders or weaken as more topics are considered. In the example shown in Figure 3a, the level of alignment observed between opinions on economic policy and immigration remains almost the same when adding a third issue (i.e., education) and even when all four topics are included. This indicates that the alignment is robust across multiple dimensions, and individuals who have a specific stance on one topic (e.g., immigration) are likely to have predictable stances on other topics (e.g., economic policy and education). Conversely, if the alignment dropped significantly when more topics are considered, the spectrum would indicate that the pairwise alignments do not extend to a broader set of issues, revealing the existence of cross-cutting cleavages.
In general, when using $A_{\text{AMI}}$, the theoretical upper bound of the alignment spectrum is represented by the horizontal line at $A=1$, indicating perfect alignment across all the issues. In contrast, the alignment values for a completely random system would be close to 00 at any order. As these bounds represent two extremes of ideological coherence, the alignment spectrum of a real-world system can describe various types of constraints within this range.
Figure 3: An example of alignment spectrum from online discussions related to four topics before Finnish elections 2023 (Panel a). The x-axis indicates the size of the topic combinations, which corresponds to the order of alignment, while the y-axis shows the corresponding alignment scores. Each point in the figure indicates a specific subset of topics, as indicated by the icons. Panel b, d, f show different fundamental patterns of higher-order alignment that we observe when analyzing real-world systems with our proposed measure. Moderate or high multiway alignment (Panel b) is usually observed in legislative systems, such as Finnish Parliament (Panel c). Panel d shows that, even when pairwise alignment is moderate, alignment at higher-order might be low. An example of this is shown in Panel e, where alignment at order $>3$ rapidly decreases towards 0. Finally, Panel f shows an example of mixed pattern of higher-order alignment, where some topics are aligned and others are not. This was the case in Finnish parliamentary discussions during 2020–2021 (Panel g). The black markers in Panel c, e, g show values of alignment that are not statistically significant, as they fall within the $95\%$ confidence interval obtained by the null model. In Panel a, all points are significant with respect to the null model.
@@ 4 Alignment Spectrum
Figure 3 illustrates two fundamental patterns that may emerge when analyzing political alignment across multiple issues using our proposed measure. In cases of moderate or high multiway alignment, individuals tend to hold consistent views across issues, leading to a funnel-shaped alignment spectrum where multiway alignment scores remain significantly above zero even when considering all topics (Figure 3b). This pattern is commonly observed in legislative systems, where the alignment across different topics reflects the dynamics of policymaking and political discourse. Figure 3c shows the alignment spectrum of Finnish Parliament in 2019–2020, which demonstrates alignment that could originate from adherence to party lines, whereby members from the same party (or coalition) vote in a similar way on a wide range of issues. The same effect is also visible when analyzing voting behavior in U.S. House (see Appendix H). Being a two-party system, the U.S. exhibits a much more pronounced effect of elite polarization on the spectrum of multiway alignment, and the multiway alignment score usually remains at moderate values even when considering as many as $18$ different topics. Yet, the pattern in Figure 3b is not limited to legislative systems. We observe a similar alignment spectrum also for social media discussions centered around political topics before Finnish elections (Figure 3a). Although the emergence of alignment within discussions around political parties is anticipated in a multi-party system, we discuss in Appendix G how the observed increase in higher-order alignment across issues signifies a shift towards a more partisan-centric political landscape.
The pattern of alignment in Figure 3b occurs because individuals maintain some level of alignment across multiple topics. At the top of the spectrum, scores are high because opinion groups tend to remain similar across the issues. As the number of issues increases (from pairs to triads and beyond), the alignment score decreases. However, when opinion groups are generally aligned, this decline is gradual, forming the funnel shape. The bottom part of the funnel represents the minimum alignment observed for each number of issues. If this minimum score is significantly above zero, it suggests the existence of constraints that could span a larger set of topics. Conversely, if the system does not exhibit higher-order alignment, the alignment scores drop close to zero as more issues are included.
This pattern of low higher-order alignment manifests as a triangular-shaped spectrum that rapidly decreases toward zero as more issues are considered (Figure 3d). The lower triangle shape indicates that individuals’ opinions across different topics are largely independent at higher-orders, or that the stances over larger sets of issues become cross-cutting. In this scenario, knowing someone’s position on one issue might provide little to no information about their stance on a combination of other issues. Figure 3e shows an example of this type of spectrum, that starts with high alignment scores for pairs of issues but rapidly declines as more issues are added.
The alignment spectrum can also exhibit a mixed pattern, where some issues are aligned while others show little to no alignment. An example of this is the distinctive double-triangle shape in Figure 3f. This mixed pattern consists of an upper funnel-like section and a lower triangular section. The upper triangle, resembling a small funnel, represents the subset of issues where individuals tend to be aligned. For these issues, alignment scores remain relatively high, even as the number of issues considered increases, but only for a subset of the topics. The lower triangle mirrors a low-alignment pattern, where scores quickly approach zero as more issues are introduced. This double-triangle pattern was observed in the Finnish Parliament during the COVID-19 pandemic (Figure 3g), and reveals a known phenomenon that would be missed in a purely pairwise analysis. Notably, during the COVID-19 pandemic, especially early on, Finnish parliamentary politics exhibited unconventionally high levels of cooperation between the government and opposition coalitions (Lehtonen and Ylä-Anttila, 2024). This was evident in how both sides repeatedly appealed for broad cooperation, reciprocally praise each other for implemented COVID responses, and even in how the opposition withdrew their interpellation on the government’s immigration policy (Eduskunta.fi, 2020). Despite this drastic shift from conventional politics, the changes brought by the pandemic become glaringly evident only when analyzing the patterns of alignment at higher orders, where the double-triangle shape of the alignment spectrum shows two clear clusters of topics, one retaining alignment levels comparable to the previous period, and the other exhibiting markedly low alignment. As discussed in Appendix I, these patterns reflect the collaborative dynamics explicitly adopted to face the crisis.
5 Case Study: Higher-Order Partisan Sorting of the U.S. Public
@@
Building on concepts developed in the previous sections, we now turn our attention to a practical application through a case study that focuses on the alignment of political attitudes in the U.S. public. By analyzing the multiway alignment scores across multiple political topics, we aim to characterize partisan sorting in the United States. Our analysis uses data from the American National Election Studies (ANES) surveys to understand how political attitudes align across various topical issues.
Established in 1948, the ANES surveys are conducted before and after every presidential election to investigate voter attitudes, party identification, and electoral behavior over time (American National Election Studies, 2021). ANES surveys have been used to study polarization, particularly affective polarization, by applying traditional statistical methods such as factor analysis (Abramowitz, 2021), Cohen’s d coefficient (Cavari and Freedman, 2018), or regression models (Enders and Lupton, 2021; Lupton et al., 2020; Robison and Moskowitz, 2019) to identify underlying dimensions of political attitudes that contribute to polarization, measure the effect size of differences in political attitudes between groups, and quantify the relationships between demographic factors and the degree of polarization among the electorate.
Traditionally, studies on public opinion polarization have focused on individual issue attitudes, often overlooking how these attitudes are interconnected across multiple issues. This approach misses the broader picture of how people’s opinions form a cohesive ideological framework and potentially sort individuals into opposing factions. As Baldassarri and Gelman (2008) argue, opinion alignment, rather than just opinion radicalization, affects social integration and political stability. Similarly, other authors have argued that political polarization is driven by social group identification and outgroup aversion (Bougher, 2017; Mason, 2018). Pursuing this idea, many authors have used ANES surveys to study how individuals’ attitudes toward various issues align with party identification or with other issues (Baldassarri and Gelman, 2008; Kozlowski and Murphy, 2021; Bougher, 2017; Hetherington, 2009; Lüders et al., 2024). Still, the study of alignment has mostly been approached by measuring pairwise correlation between pairs of issues or between each issue and party identification (Baldassarri and Gelman, 2008; Kozlowski and Murphy, 2021; Hetherington, 2009).
An attempt to consider multiple issues at the same time can be found in Bougher (2017), where partisan alignment is measured by considering whether individuals’ stances on an issue align with the opposing party’s ideological position, with the own party’s position, or with neither one. Lüders et al. (2024) adopt a network-based approach to study attitude alignment over eight issues, constructing and visualizing a graph where nodes are stances and edges are weighted based on phi correlation coefficients between each pair of stances. While these studies focus on the organization of public opinion around ideologically- or partisan-opposed factions that contain multiple attitudes, neither explicitly examine higher-order dependencies among the attitudes.
In this case study, we aim to characterize partisan-ideological sorting in terms of the topical issues which ideological groups coalesce their identities around. To do so, we analyze multiway alignment (both issue alignment and partisan alignment) in U.S. public opinion using the ANES Time Series data from 2004–2020. Instead of directly measuring polarization, we examine how the alignment of opinions across various issues gives rise to ideological groups and partisanship.
@@
Our analysis reveals several key findings. First, we observe an increase in multiway alignment over time, indicating that the interdependencies among the topics in analysis (seven issues and partisan identification) have intensified. This trend is shown in Figure 4, where we compare the full spectrum of the multiway alignment computed from the 2004 and 2020 data. While Figure 4 suggests a more polarized landscape in contemporary American discourse, we also find that the alignment of stances concerning abortion and LGBT rights, which was predominant in 2004, has been replaced by a more pronounced alignment between gun access and immigration in 2020. Notably, the alignment score pertaining to the latter pair surpasses the three highest scores observed in 2004, signifying the existence of stronger 2-way constraints and the changing salience of different issues over time. Moreover, our results show that the overall multiway alignment across all eight topics has increased, portraying a higher level of attitude constraint in the public opinion.
@@
The second key finding concerns partisan sorting. The analysis of 3-tuples and higher illustrates the growing importance of political partisanship in U.S. public opinion. Whereas in 2004, positions toward immigration played a central role within triadic-and-higher relationships, in 2020, party preferences became one of the most prominent topics, marking a shift towards sociopolitical affiliations as key alignment determinants. In this respect, our measure of multiway alignment allows us to investigate research questions that study partisanship across a multitude of issues. Specifically, we are now able to quantify the strength of the relationship between party preference and a bundle of multiple topics.